The firm's long-run total cost is given by LTC = 5,000Q - 100Q^2 + Q^3 and its long-run marginal cost is given by LMC = 5,000 - 200Q + 3Q^2. At what output level does the firm experience diseconomies of scale?

I don't understand how it goes from this " 5000 - 100Q + Q2 " to this equation "-100 + 2Q = 0"
Answer #1

LTC = LATC 5000 d - lood²+ a3 LTC o - 5000 0-100 0²7 83 LATC - Sooo - loo 0 + 0²The long run average total cost, LATC es minimum when LATC = LMC. LMC= 5000 - 2000+ 302 So, LATC= Lmc 50oo-lood to2 = 5000- 2